### v1

# Ultrastable 2D glasses and packings explained by local centrosymmetry

Abstract
Using the most recent numerical data by Bolton-Lum et al. [Phys. Rev. Lett. 136, 058201 (2026)], we demonstrate that ideal ultrastable glasses in the athermal limit (or ultrastable ideal 2D disk packings) possess a remarkably high degree of local centrosymmetry. In particular, we find that the inversion-symmetry order parameter for local force transmission introduced in Milkus and Zaccone, [Phys. Rev. 93, 094204 (2016)], is as high as $F_{IS}=0.93546$, to be compared with $F_{IS}=1$ for perfect centrosymmetric crystals free of defects, and with $F_{IS}\sim 0.3-0.5$ for standard random packings. This observation provides a clear, natural explanation for the ultra-high shear modulus of ideal packings and ideal glasses, because the high centrosymmetry prevents non-affine relaxations which decrease the shear modulus. The same mechanism explains the absence of boson peak-like soft vibrational modes. These results also confirm what was found previous work, i.e. that the bond-orientational order parameter is a very poor correlator for the vibrational and mechanical properties of disordered packings.
Ideal or ultrastable glasses have emerged as a useful groundwork to study long-standing open issues such as the nature of the glass transition. Understanding their properties also bears relevance for experimental systems, where ultrastable glasses have been realized since their discovery by Ediger and co-workers in 2007 using vapor deposition techniques [17]. Yet, there are still a lot of open questions, and often misconceptions, about the origin of the unusual properties of ultrastable glasses, which appear to be more similar to those of crystals [7, 21]. It thus appears like a paradox, that these systems have mechanical and vibrational properties akin to crystals, but their structure is completely disordered and random, as reflected in low values of traditional metrics such as the bond-orientational order parameter.
In this paper, we provide the missing geometric link without which it is impossible to resolve the above paradox. By analyzing the ultrastable 2D random disk packings of Ref. [3], we compute a metric which captures the local statistical degree of inversion-symmetry around each disk. This metric directly connects to the microscopic origin of nonaffine displacements in the deformation and lattice dynamics of amorphous materials. The core concept is easily explained: in a perfect centrosymmetric crystal subjected to a shear deformation, each particle will receive forces from its nearest-neighbours which cancel each other out by inversion-symmetry. Hence, each particle is already at mechanical equilibrium in the (affine) position prescribed by the deformation tensor. Conversely, in a disordered lattice with no centrosymmetry, under strain, each particle will receive forces from its nearest-neighbors, which do not cancel each other out by symmetry, thus leaving a net unbalanced force in the affine position [22, 6, 13, 25, 24]. This net force, in turn, triggers an additional (nonaffine) displacement, on top of the affine one dictated by the applied strain. The nonaffine displacements occur in the local force field of interparticle interactions and, because force times displacement is a work, specifically an internal work done by the particles to keep mechanical equilibrium, this contributes negatively to the internal energy of deformation and leads to a negative contribution to the shear modulus. The mechanism is schematically illustrated in Fig. 1.
Figure 1: Schematic illustration of the connection between local inversion symmetry and nonaffine displacements in disordered solids. Top: An imposed shear deformation defines an affine displacement field (dashed lines). In inversion–asymmetric environments particles experience a net force and undergo nonaffine displacements away from the dashed lines. Bottom: In centrosymmetric environments the sum of forces vanishes, while inversion-symmetry breaking induces a finite force imbalance in the affine position that drives nonaffine rearrangements. This force–imbalance mechanism underlies nonaffinity in disordered solids [23, 6, 22].
For isostatic packings, it has been shown rigorously [23] that this negative nonaffine contribution exactly cancels the affine contribution, thus leading to the vanishing of rigidity or unjamming at $z_{c}=2d$ for central-force packings at the random close packing (RCP) [26], where $z$ is the coordination number and $d$ the space dimension. Therefore, the statistical degree of centrosymmetry in local force-transmission, introduced in Ref. [11], can be leveraged as a useful metric to predict the mechanical and vibrational properties of a amorphous systems solely based on the static particle-level structure of the system, and this usefulness has been benchmarked several times in the past on experimental data for very different systems such as colloidal glasses [9, 10], jammed packings [20] and metallic glasses [4]. Furthermore, it was shown already in [11] that the inversion-symmetry order parameter $F_{IS}$ not only provides a strong correlation with the mechanical response but also with the vibrational properties, such as the boson peak excess of soft vibrational modes over the Debye prediction. At the same time, a traditional metric such as the bond-orientational order parameter [15] is unable to correlate with the mechanical and vibrational properties of disordered systems such as random packings and random networks [11].
Bolton–Lum et al. [3] recently introduced numerically generated two-dimensional polydisperse disk packings that exhibit several striking properties, including hyperuniformity [19], a finite shear modulus in the low-pressure limit, and a vibrational spectrum lacking boson-peak–like excess modes. The latter two features are unusual for standard disordered packings [12, 18] and suggest a predominantly affine elastic response. Furthermore, the packing fraction found in [3] of about $0.91$ is significantly lower than the maximal packing fractions for random disk packing of polydisperse disks which may exceed $0.96$ for very wide size distributions [27]. The structural origin of these crystal-like properties, however, remains unclear.
We now test this hypothesis directly on the numerical data of Bolton-Lum et al. [3]. In particular, we analyze the ideal disk packing sample shown in Fig. 2, and we compute the inversion-symmetry order parameter $F_{IS}$ as defined in [11].
Figure 2: (a) Numerical packing of polydisperse disks. Panel (a) is reproduced from Bolton–Lum et al., [3], with permission from the American Physical Society. (b) Gabriel-graph constructed from the particle centers corresponding to the configuration in (a), used to define nearest-neighbor bonds for the computation of the inversion-symmetry order parameter $F_{\mathrm{IS}}$.
With reference to Fig. 1, in the nonaffine lattice-dynamics framework of Milkus and Zaccone [11], the microscopic driver of nonaffine motion is the force imbalance that appears when atoms are brought to their affine positions under an imposed strain: in a locally centrosymmetric environment the nearest-neighbor forces cancel, whereas local inversion-symmetry breaking generically produces a nonzero net force that must be released by additional nonaffine displacements. In the harmonic approximation this imbalance can be written as $\mathbf{f}_{i}=\boldsymbol{\Xi}_{i}\,\gamma$, where $\gamma$ is the shear strain, and the vector $\boldsymbol{\Xi}_{i}$ defined as [6, 22]:
$\boldsymbol{\Xi}_{i}=-\kappa\sum_{j}R_{ij}\mathbf{n}_{ij}\,n_{ij}^{x}\,n_{ij}^{y}$ (1)
where $\kappa$ is the spring constant of the particle-particle interaction, $R_{ij}$ is the scalar distance between particle $i$ and its nearest-neighbor $j$, and $\mathbf{n}_{ij}$ is the unit vector from $i$ to $j$. The affine force vector $\boldsymbol{\Xi}_{i}$ encodes the degree of local inversion-symmetry breaking at site $i$ (it vanishes identically for locally centrosymmetric environments and is finite otherwise), thereby directly controlling the magnitude of nonaffine relaxations and the associated nonaffine softening of the elastic response. To quantify inversion-symmetry breaking as a structural indicator of shear rigidity, they define a normalized order parameter $F_{IS}$, based on the total affine force field magnitude $|\boldsymbol{\Xi}|^{2}$, namely
$F_{\mathrm{IS}}\;=\;1-\frac{|\boldsymbol{\Xi}|^{2}}{\kappa^{2}\sum_{ij}\!R_{ij}^{2}\left(n^{x}_{ij}n^{y}_{ij}\right)^{2}},$ (2)
which equals unity for perfect centrosymmetric lattices (where $|\boldsymbol{\Xi}|^{2}=0$) and tends to zero in the limit of maximally broken inversion symmetry, thus providing a direct structural predictor for the onset of nonaffine motions and the accompanying anomalies in shear elasticity and vibrational spectra. In the above formula, [11, 22]:
$|\boldsymbol{\Xi}|^{2}=\kappa^{2}R_{0}^{2}\sum_{i}\sum_{\alpha\in\{x,y\}}\left(\sum_{j\in\mathrm{nn}(i)}n_{ij}^{\alpha}\,n_{ij}^{x}\,n_{ij}^{y}\right)^{2}.$ (3)
To identify nearest-neighbor pairs for the evaluation of the inversion-symmetry order parameter $F_{\mathrm{IS}}$, we first constructed the Delaunay triangulation of the particle centers from Fig. 2(a) retained within the observation window. The corresponding Gabriel graph, shown in Fig. 2(b) was then obtained by retaining only those Delaunay edges for which the closed disk having the edge as diameter contains no other particle centers. This procedure yields a parameter-free definition of nearest neighbors that is robust to local density fluctuations and does not rely on an arbitrary distance cutoff. The order parameter $F_{\mathrm{IS}}$ was subsequently evaluated as a bond-averaged quantity over the resulting set of Gabriel-connected particle pairs, using the pairwise separation $R_{ij}$ for each contributing bond.
From the numerical configuration comprising $288$ disk centers in Fig. 2(a), particles whose centers lay outside the cropped observation window were excluded from the analysis. Specifically, a particle $i$ with coordinates $(x_{i},y_{i})$ was retained only if $0\leq x_{i}\leq 1$ and $0\leq y_{i}\leq 1$; all others were discarded. This criterion removed $33$ particles located near the boundaries, corresponding to an excluded fraction of approximately $11.5\%$ of the total population, leaving $N=255$ particles for the computation of structural observables. No additional distance-based or neighbor-based buffer was imposed: the exclusion was determined solely by the particle-center coordinates. The inversion-symmetry order parameter $F_{\mathrm{IS}}$ was then evaluated using the resulting bulk configuration, ensuring that all contributing interparticle bonds were fully contained within the observation window and free of boundary-induced incompleteness.
From the computation, we finally obtain the following estimate of $F_{IS}$ for the ultrastable ideal disk packing of Bolton-Lum et al. [3]:
$F_{IS}=0.93546$ (4)
which is indeed very close to the ideal defect-free centrosymmetric crystal limit, $F_{IS}=1$ [11].
For completeness, we also compute the centrosymmetry parameter of Kelchner, Plimpton, and Hamilton [5], defined as:
$CS=\sum_{i=1}^{N/2}\left|\mathbf{R}_{i}+\mathbf{R}_{i+N/2}\right|^{2}.$ (5)
Here the $N$ nearest neighbors of a given particle are first identified, and $\mathbf{R}_{i}$ and $\mathbf{R}_{i+N/2}$ denote the vectors from the central particle to a particular pair of nearest neighbors. Among the N(N1)/2N(N-1)/2 possible neighbor pairs, the quantity $\lvert\mathbf{R}_{i}+\mathbf{R}_{j}\rvert^{2}$ is evaluated for each pair, and the $N/2$ smallest values are selected for the sum. These pairs typically correspond to neighbors located in approximately opposite directions with respect to the central particle, motivating the $i+N/2$ notation. The parameter $N$ is a user-defined even integer chosen to match the coordination number of the local environment. In a perfectly centrosymmetric environment, opposite neighbor vectors cancel pairwise and $CS=0$, while local distortions and defects lead to finite values of $CS$.
Using the same Gabriel graph employed for the computation of the inversion-symmetry order parameter $F_{\mathrm{IS}}$ above, we also evaluated the centrosymmetry parameter $CS$ for each particle. Nearest neighbors were identified from the Gabriel-graph connectivity, and the centrosymmetry parameter was computed using $N=6$, appropriate for two-dimensional ideal disk packings. Averaging over all particles within the bulk region, we obtain a mean value $\langle CS\rangle\simeq 4.6\times 10^{-3}$, with a median of comparable magnitude. To facilitate comparison with previous work, we also consider a normalized centrosymmetry parameter $CS^{\ast}=CS/R_{0}^{2}$, where $R_{0}$ is the mean nearest-neighbor distance. Using the same Gabriel-defined neighbor network and bulk particle set as for the inversion-symmetry order parameter $F_{\mathrm{IS}}$, we find $\langle CS\rangle\simeq 4.6\times 10^{-3}$ and $R_{0}\simeq 7.1\times 10^{-2}$, yielding a normalized value $\langle CS^{\ast}\rangle\simeq 0.92$. This magnitude is comparable to values reported for weakly distorted crystalline environments and is significantly smaller than those associated with strong defects such as dislocation cores or free surfaces. The result is consistent with the large value of $F_{\mathrm{IS}}$ obtained for the same configuration above, indicating that local environments remain close to centrosymmetric despite geometric disorder.
In particular, the magnitude of the normalized centrosymmetry parameter obtained here, $\langle CS^{\ast}\rangle\simeq 0.9$, is comparable to values reported for weakly distorted crystalline environments subject to thermal fluctuations or elastic strain, for which $CS^{\ast}\sim 0.1{-}1$ [5, 16, 14]. These values are significantly smaller than those associated with stacking faults or twin boundaries ($CS^{\ast}\sim 2{-}6$) and are more than an order of magnitude smaller than those characteristic of dislocation cores or free surfaces ($CS^{\ast}\gtrsim 5$) [5, 8]. The observed high centrosymmetry is therefore consistent with locally crystalline-like environments exhibiting geometric disorder but lacking strong topological-like defects [1, 2].
The central result of this work is the demonstration that ideal two-dimensional disk packings and ultrastable glasses possess an exceptionally high degree of local inversion symmetry, quantitatively captured by the inversion-symmetry order parameter $F_{\mathrm{IS}}$. The measured value $F_{\mathrm{IS}}=0.93546$ is remarkably close to the crystalline limit $F_{\mathrm{IS}}=1$ and stands in sharp contrast to the much smaller values ($F_{\mathrm{IS}}\sim 0.3-0.5$) characteristic of ordinary jammed packings and structural glasses. This finding provides a direct and physically transparent structural explanation for the long-standing puzzle of why ideal glasses exhibit crystal-like mechanical and vibrational properties despite the absence of long-range order. The large value of $F_{\mathrm{IS}}$ measured here naturally accounts for both the anomalously high rigidity and the crystal-like vibrational spectrum reported for ideal disk packings, without invoking hidden crystalline order or exotic mechanisms.
The present analysis also clarifies the limitations of traditional structural metrics. Bond-orientational order parameters, while sensitive to angular correlations, are fundamentally blind to inversion symmetry and therefore fail to correlate with mechanical stiffness or vibrational anomalies in disordered solids. In contrast, $F_{\mathrm{IS}}$ directly probes the symmetry properties of local force transmission and thus provides a robust structural predictor of elastic and vibrational behavior. The consistency between the large $F_{\mathrm{IS}}$ and the small, crystal-like value of the normalized centrosymmetry parameter $CS^{\ast}$ further reinforces this interpretation.
More broadly, these results suggest that local centrosymmetry constitutes a key organizing principle for amorphous solids at the edge between glassy and crystalline behavior. Ideal glasses and ideal packings emerge as systems that remain topologically disordered yet are statistically optimized to maximize local inversion symmetry, thereby suppressing nonaffine softening. This insight offers a unifying geometric framework to understand ultrastability, hyperuniform mechanical response, and the disappearance of low-frequency vibrational excitations. Beyond two-dimensional disk packings, the same concepts are expected to apply to three-dimensional glasses, jammed materials, and metallic glasses, opening the door to symmetry-based design principles for mechanically robust amorphous materials.
Acknowledgments
A.Z. gratefully acknowledges funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”, from US Army Research Office through contract nr. W911NF-22-2-0256, and from the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program.